OpenStudy (idealist10):

On a multiple-choice exam with 3 possible answers for each of the 5 questions, find the probability that a student would get 4 or more correct answers just by guessing.

OpenStudy (idealist10):

@ganeshie8 @wio @radar

ganeshie8 (ganeshie8):

binomial distribution ?

OpenStudy (idealist10):

I think so.

ganeshie8 (ganeshie8):

just addup the probabilities p(X>=4) = P(4) + P(5)

ganeshie8 (ganeshie8):

start by finding the probabiltiy for sucess in guessing a correct answer

OpenStudy (idealist10):

So is it P(4)+P(5)=(1/3)^4+(1/3)^5?

ganeshie8 (ganeshie8):

thats a good try, but no P(5) equals (1/3)^5 P(4) is bit tricky as you have a choice to tick false option for any one of the five questions

ganeshie8 (ganeshie8):

S : success F : failure

ganeshie8 (ganeshie8):

you can answer exactly four correct answers in any one of below 5 ways ``` FSSSS SFSSS SSFSS SSSFS SSSSF ``` yes ?

OpenStudy (idealist10):

Yes. And then?

ganeshie8 (ganeshie8):

Notice each of them have four success ticks and one failure tick so the probability of each of them is same : (1/3)^4 * (2/3)^1

ganeshie8 (ganeshie8):

P(4) = 5 * (1/3)^4 * (2/3)^1

ganeshie8 (ganeshie8):

see if that makes sense more or less ^

OpenStudy (idealist10):

Yes, it makes sense! Thanks!

ganeshie8 (ganeshie8):

yw :) add that to P(5) : P(X>=4) = P(4) + P(5) = 5 * (1/3)^4 * (2/3)^1 + (1/3)^5

OpenStudy (idealist10):

I did. :)